# Spatio-temporal point process intensity estimation using zero-deflated subsampling applied to a lightning strikes dataset in France

### v1

@@ Abstract
[…]
@@ 1 Introduction
Severe thunderstorms can be associated with personal injury or costly infrastructure damage. Among the major risks to be assessed are those associated with high lightning intensity, and they have been studied for about twenty years (see e.g. Curran et al. (2000), Mona et al. (2016), Schulz et al. (2005, 2016), Simon et al. (2017), Hernandez-Magallanes and Genton (2019), Nampak et al. (2021)). Additionally, the interest in evaluating the intensity of cloud-to-ground lightning strikes might increase in the context of an energy transition: wind farms are multiplying, and the risk of lightning is one of the five hazards typically identified by manufacturers. From a physical standpoint, a thunderstorm corresponds to the dissipation of energy involving various thermodynamic processes, such as convection. To describe these complex processes, several covariates are known to be useful: the CAPE (Convective Available Potential Energy), which acts as the ’fuel’ in the development of the cloud; the $\theta_{w}^{\prime}$ at 850 hPa (Wet-bulb potential temperature at 850 hPa), which is known to be a good synoptic predictor, giving important information on the position of the fronts; the temperature at two different altitudes (20m and 1500m), providing information on the vertical profile of the air mass (convection); humidity; and wind (zonal and meridional components). A few other purely spatial covariates like the altitude, the distance to the sea should also have an influence (see Section 2 for more details). Lightning strike impacts, although continuously studied by scientists in the fields of physics, climatology or statistics remains a phenomenon with an important part of randomness. The set of strikes collected over time and space constitute a spatio-temporal point pattern and the goal of this paper is to characterize its distribution in terms of climatological, topographical covariates.
@@ 2 Presentation of data and their complexity
[…]
@@ 5.2 Subsampled version of $\mathrm{BRL}_{\text{\scriptsize\rm logit}}$
Now, let us keep the spirit of the method $\mathrm{BRL}_{\text{\scriptsize\rm logit}}$ and (wrongly) assume that $\mathrm{logit}\mathbb{P}(I_{j}=1)=\log(\delta_{j})+\boldsymbol{\beta}^{\top}\mathbf{C}_{j}$. Then, Equation (5.1) becomes the cornerstone of Section 5: the likelihood of the $I_{j}\mid j\in\mathcal{J}$ can still be used to estimate $\boldsymbol{\beta}$ using a standard logistic regression, but now with offset term $(\log(\delta_{j})+\log(\pi_{1}/\pi_{0}))_{j\in\mathcal{J}}$.
@@ 5.4 Subsampled version of $\mathrm{CLRL}$ and $\mathrm{WCLRL}$
Following Sections 4.2 and 5.2, we can propose the following subsampled version of $\mathrm{CLRL}$ (the subsampled version of $\mathrm{WCLRL}$ is similar). Let $\mathcal{N}$ (resp. $\mathcal{J}$) be a sample of $\{j=1,\dots,N+J:\tilde{I}_{j}=1\}$ (resp. $\{j=1,\dots,N+J:\tilde{I}_{j}=0\}$) with inclusion probabilities $\pi_{1}=\mathbb{P}(j\in\mathcal{N}\mid\tilde{I}_{j}=1)$, $\pi_{0}=\mathbb{P}(j\in\mathcal{J}\mid\tilde{I}_{j}=0)$ and let $\mathcal{K}=\mathcal{N}\cup\mathcal{J}$. Then, it is clear from Section 5.2 that
@@ 5.5 Subsampled version of $\mathrm{PL}$
Assuming $\tau_{0}=\pi_{0}$ and $\tau_{\ell}=\pi_{1}$ for any $\ell\geq 1$, which implies we sample according to the values of $I_{j}$ as in Sections 5.2-5.3, leads to
@@ 5.6 A brief simulation study, remarks and possible extensions
Method
${\mathrm{BRL}_{\text{\scriptsize\rm logit}}}$
$\mathrm{BRL}_{\text{\scriptsize\rm cloglog}}$
$\mathrm{CLRL}$
$\mathrm{WCLRL}$
$\mathrm{PL}$
(see Section 4.3)
(see Section 4.4)
(see Section 4.2)
(see Section 4.2)
(see Section 4.1)
GLM family
binomial binomial binomial binomial poisson
Link
logit cloglog logit logit log
R formula
I~C I~C Icheck~Ccheck
Itilde~Ctilde, w=wtilde
N~C
Offset term
$(\log\delta_{j})_{j}$
$(\log\delta_{j})_{j}$
$(\log\delta_{j})_{j}$
$(\log\delta_{j})_{j}$
$(\log\delta_{j})_{j}$
Size of the response vector
$J$
$J$
$N+J$
#$\#J_{1}+J$
Unbiased score
$\times$
$\checkmark$ (if Poisson)
$\checkmark$
$\checkmark$
$\checkmark$
Robustness
$\times$
$\times$
$\checkmark$
$\checkmark$
$\checkmark$
Variance
Table 1: Summary of subsampled versions of parametric methods used to estimate an exponential family model with piecewise constant covariates; I (resp. Icheck,Itilde,N) stands for the vector of the $I_{j}$ (resp. $\check{I}_{j}$, $\tilde{I}_{j}$, $N_{j}$); C (resp. Ccheck,Ctilde) stands for the matrix of constant covariates; $\delta_{j}$ is the volume of the spatio-temporal cell $\Delta_{j}$.
Subsampled version of
${\mathrm{BRL}_{\text{\scriptsize\rm logit}}}$
$\mathrm{BRL}_{\text{\scriptsize\rm cloglog}}$
$\mathrm{CLRL}$
$\mathrm{WCLRL}$
$\mathrm{PL}$
(see Section 5.2)
(see Section 5.3)
(see Section 5.4)
(see Section 5.4)
(see Section 5.5)
GLM family
binomial binomial binomial binomial poisson
Link
logit cloglog.sub logit logit log
R formula
I.sub~ C.sub
I.sub~ C.sub
Icheck.sub~ Ccheck.sub
Itilde.sub~ Ctilde.sub, w=wtilde.sub
N.sub~ C.sub
Offset term
$\big{(}\log\frac{\pi_{1}}{\pi_{0}}$ $+\log\delta_{j}\big{)}_{j}$
$(\log\delta_{j})_{j}$
$\big{(}\log\frac{\pi_{1}}{\pi_{0}}$ $+\log\delta_{j}\big{)}_{j}$
$\big{(}\log\frac{\pi_{1}}{\pi_{0}}$ $+\log\delta_{j}\big{)}_{j}$
$\big{(}\log\big{(}\frac{\pi_{1}}{\pi_{0}}\big{)}$ $\mathbf{1}(j\in\mathcal{J}_{0})$ $+\log\left(\delta_{j}\right)\big{)}_{j}$
Size of the response vector
#$\#\mathcal{J}_{0}+\#\mathcal{J}_{1}$
#$\#\mathcal{J}_{0}+\#\mathcal{J}_{1}$
#$\#\mathcal{N}+\#\mathcal{J}$
#$\#\mathcal{N}_{1}+\#\mathcal{J}$
#$\#\mathcal{J}_{0}+\#\mathcal{J}_{1}$
Unbiased score given subsample
$\times$
$\checkmark$ (if Poisson)
$\checkmark$
$\checkmark$
@@
Robustness
$\times$
$\times$
$\checkmark$
$\checkmark$
$\checkmark$
Table 2: Summary of characteristics of parametric methods used to estimate an exponential family model with piecewise constant covariates; I.sub (resp. Itilde.sub,Icheck.sub,N.sub,), stands for a (dependent) subsample of the $I_{j}$ (resp. $\tilde{I}_{j}$, $\check{I}_{j}$, $N_{j}$); C.sub (resp. Ccheck.sub,Ctilde.sub) stands for the subsampled matrix of constant covariates; $\delta_{j}$ is the volume of the spatio-temporal cell $\Delta_{j}$; (see Section 5 for a definition $\pi_{0}$, $\pi_{1}$, $\mathcal{J}_{0}$, $\mathcal{J}_{1}$, $\mathcal{N}$, $\mathcal{J}$ and $\mathcal{N}_{1}$ according to each method).
@@ 6.2 Results
Tables 3-4 summarize the results obtained. We report the values of prediction critera PPV, AUC and WW detailed in Section 6.1, as well as the calculation time per bag. From these two tables, a few comments can be formulated: first of all, whatever the criterion, it seems useless to choose more that one bag; second, note that AUC does not appear so informative, since all methods and subsampling rates achieve the same kind of performance. This is also the case for the PPV criterion in winter 00:00-06:00 (Table 4). Note also that PPV is quite small, in particular for winter (see Table 4). This confirms that the prediction of lightning storms is a very difficult task. However, this should be compared to the PPV achieved by the random classifier, which is the proportion of non-empty cells for these pairs of (season, moment) ($0.0154$ for (summer, 12:00-18:00) and $0.0004$ for (winter, 00:00-06:00)). Two methods ($\mathrm{PL}$ and $\mathrm{BRL}_{\text{\scriptsize\rm cloglog}}$) stand out for their stability concerning the subsampling rate, as demonstrated by the behaviors of PPV and WW in Table 3. Additionally, in the absence of subsampling ($\pi_{0}=1$), PPV lacks discrimination. However, in terms of WW, $\mathrm{PL}$ and $\mathrm{WCLRL}$ slightly outperform the other methods. Notably, subsampling significantly impacts computational time reduction (by at least a factor of 40) without compromising results.
@@ 7 Conclusion
We believe this work leads to several perspectives. It would be first interesting to study more deeply these procedures. For instance, can we still be ensured that the estimator derived by these procedures is consistent and asymptotically normal? Extending these procedures for non piecewise constant covariates is a second perspective. Finally, investigating inhomogeneous subsampling, i.e. for instance letting $\pi_{0}$ and $\pi_{1}$ depend on the cell $\Delta_{j}$ is also a direction for further research.
Regarding the dataset application, modelling second-order characteristics of the process or even better modelling the distribution of $\mathbf{Y}$ for instance as a mixture of a temporal Hawkes process and Neymann-Scott or log-Gaussian spatial Cox process for the spatial part could be an interesting first attempt. Such a modelling could be pertinent if one in addition intend to provide a short-term forecast of electrical activity.

### v2

@@ Abstract
Cloud-to-ground lightning strikes observed in a specific geographical domain over time can be naturally modeled by a spatio-temporal point process. Our focus lies in the parametric estimation of its intensity function, incorporating both spatial factors (such as altitude) and spatio-temporal covariates (such as field temperature, precipitation, etc.). The events are observed in France over a span of three years. Spatio-temporal covariates are observed with resolution $0.1^{\circ}\times 0.1^{\circ}$ ($\approx 100$km2) and six-hour periods. This results in an extensive dataset, further characterized by a significant excess of zeroes (i.e., spatio-temporal cells with no observed events). We reexamine composite likelihood methods commonly employed for spatial point processes, especially in situations where covariates are piecewise constant. Additionally, we extend these methods to account for zero-deflated subsampling, a strategy involving dependent subsampling, with a focus on selecting more cells in regions where events are observed. A simulation study is conducted to illustrate these novel methodologies, followed by their application to the dataset of lightning strikes.
@@ 1 Introduction
Severe thunderstorms can be associated with personal injury or costly infrastructure damage. Among the major risks to be assessed are those associated with high lightning intensity, and they have been studied for about twenty years (see e.g. Curran et al. (2000), Mona et al. (2016), Schulz et al. (2005, 2016), Simon et al. (2017), Hernandez-Magallanes and Genton (2019), Nampak et al. (2021)). Additionally, the interest in evaluating the intensity of cloud-to-ground lightning strikes might increase in the context of an energy transition: wind farms are multiplying, and the risk of lightning is one of the five hazards typically identified by manufacturers. From a physical standpoint, a thunderstorm corresponds to the dissipation of energy involving various thermodynamic processes, such as convection. To describe these complex processes, several covariates are known to be useful (e.g. Bertram and Mayr, 2004, Chaudhuri and Middey, 2012, Taillardat and Mestre, 2020): the CAPE (Convective Available Potential Energy), which acts as the ’fuel’ in the development of the cloud; the $\theta_{w}^{\prime}$ at 850 hPa (Wet-bulb potential temperature at 850 hPa), which is known to be a good synoptic predictor, giving important information on the position of the fronts; the temperature at two different altitudes (20m and 1500m), providing information on the vertical profile of the air mass (convection); humidity; and wind (zonal and meridional components). A few other purely spatial covariates like the altitude, the distance to the sea should also have an influence (see Section 2 for more details). Lightning strike impacts, although continuously studied by scientists in the fields of physics, climatology or statistics remains a phenomenon with an important part of randomness. The set of strikes collected over time and space constitute a spatio-temporal point pattern and the goal of this paper is to characterize its distribution in terms of climatological, topographical covariates.
@@ 2 Presentation of data and their complexity
In addition to this spatio-temporal point pattern, we also have at our disposal spatial covariates and spatio-temporal covariates used to explain the distribution of lightning strikes. Spatial covariates are observed on a regular grid of $0.1^{\circ}\times 0.1^{\circ}$ ($\approx 100$km2). They are constituted by the altitude map, the distance to the sea, as well as four additional variables computed from the AURELHY method (Bénichou, 1994). Their interest has been shown recently, see e.g. Taillardat and Mestre (2020), for statistical post-processing of ensemble forecasts. Spatio-temporal covariates consist of model outputs (Météo-France ARPEGE [^4] See e.g. https://www.umr-cnrm.fr/spip.php?article121&lang=en Model output) of climate covariates that are available only from 2013 to 2015. We have access to the following covariates observed on a regular grid of $0.1^{\circ}\times 0.1^{\circ}\times 6\text{h}$: CAPE (Convective Available Potential Energy); Humidity (at altitude 1500m); Temperatures (at altitudes 20m and 1500m); $\theta_{w}^{\prime}$ at 850 hPa (Wet-bulb potential temperature at 850 hPa); zonal and meridional components of wind. The relevance of these latter covariates has been detailed in the Introduction. The altitude map is illustrated in Figure 1. Similarly to the point pattern, these covariates obviously vary a lot, as shown in Figure 8. Figure 9 explores the interest of some of these covariates to explain the presence/absence of a lightning strike at a specific location. Maps presented in this paper do not take into consideration the earth curvature (a problem frequently pointed out in the literature Jeong and Jun (e.g. 2015)), in particular because spatial and spatio-temporal covariates provided by Météo-France already consist on grid-projected data.
@@ 5.1 Subsampled version of $\mathrm{BRL}_{\text{\scriptsize\rm logit}}$
Now, let us keep the spirit of the method $\mathrm{BRL}_{\text{\scriptsize\rm logit}}$. If
${\mathrm{logit}\mathbb{P}(I_{j}=1)=\log(\delta_{j})+\boldsymbol{\beta}^{\top}\mathbf{C}_{j},}$ (5.2)
then, the likelihood of the $I_{j}\mid j\in\mathcal{J}$ can still be used to estimate $\boldsymbol{\beta}$ using a standard logistic regression, but now with offset term $(\log(\delta_{j})+\log(\pi_{1}/\pi_{0}))_{j\in\mathcal{J}}$. As already mentioned in Section 4.3, (5.2) has no reason to hold. However, Equation (5.1) is the cornerstone of Section 5 that we exploit to provide extension of the methods $\mathrm{BRL}_{\text{\scriptsize\rm cloglog}},\mathrm{CLRL},\mathrm{WCLRL}$ and $\mathrm{PL}$.
@@ 5.3 Subsampled version of $\mathrm{CLRL}$ and $\mathrm{WCLRL}$
Following Sections 4.2 and 5.1, we can propose the following subsampled version of $\mathrm{CLRL}$ (the subsampled version of $\mathrm{WCLRL}$ is similar). Let $\mathcal{N}$ (resp. $\mathcal{J}$) be a sample of $\{j=1,\dots,N+J:\tilde{I}_{j}=1\}$ (resp. $\{j=1,\dots,N+J:\tilde{I}_{j}=0\}$) with inclusion probabilities $\pi_{1}=\mathbb{P}(j\in\mathcal{N}\mid\tilde{I}_{j}=1)$, $\pi_{0}=\mathbb{P}(j\in\mathcal{J}\mid\tilde{I}_{j}=0)$ and let $\mathcal{K}=\mathcal{N}\cup\mathcal{J}$. Then, it is clear from Section 5.1 that
@@ 5.4 Subsampled version of $\mathrm{PL}$
Assuming $\tau_{0}=\pi_{0}$ and $\tau_{\ell}=\pi_{1}$ for any $\ell\geq 1$, which implies we sample according to the values of $I_{j}$ as in Sections 5.1-5.2, leads to
@@ 5.5 A brief simulation study, remarks and possible extensions
Method ${\mathrm{BRL}_{\text{\scriptsize\rm logit}}}$ $\mathrm{BRL}_{\text{\scriptsize\rm cloglog}}$ $\mathrm{CLRL}$ $\mathrm{WCLRL}$ $\mathrm{PL}$
(see Section 4.3) (see Section 4.4) (see Section 4.2) (see Section 4.2) (see Section 4.1)
GLM family binomial binomial binomial binomial poisson
Link logit cloglog logit logit log
R formula I~C I~C Icheck~Ccheck Itilde~Ctilde, w=wtilde N~C
Offset term $(\log\delta_{j})_{j}$ $(\log\delta_{j})_{j}$ $(\log\delta_{j})_{j}$ $(\log\delta_{j})_{j}$ $(\log\delta_{j})_{j}$
Size of the response vector $J$ $J$ $N+J$$\#J_{1}+J$
Unbiased score $\times$ $\checkmark$ (if Poisson) $\checkmark$ $\checkmark$ $\checkmark$
Robustness $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Variance +/- +/- + + ++
Table 1: Summary of subsampled versions of parametric methods used to estimate an exponential family model with piecewise constant covariates; I (resp. Icheck,Itilde,N) stands for the vector of the $I_{j}$ (resp. $\check{I}_{j}$, $\tilde{I}_{j}$, $N_{j}$); C (resp. Ccheck,Ctilde) stands for the matrix of constant covariates; $\delta_{j}$ is the volume of the spatio-temporal cell $\Delta_{j}$. A $\checkmark$ (resp. $\times$) means that the statement is correct (resp. wrong). A $+$ means that the variance is shown to converge to 0 when the amount of data increases under general models. The ++ means that the variance is the lowest among considered methods. A +/- means that the variance decreases only for some specific situations.
Subsampled version of ${\mathrm{BRL}_{\text{\scriptsize\rm logit}}}$ $\mathrm{BRL}_{\text{\scriptsize\rm cloglog}}$ $\mathrm{CLRL}$ $\mathrm{WCLRL}$ $\mathrm{PL}$
(see Section 5.1) (see Section 5.2) (see Section 5.3) (see Section 5.3) (see Section 5.4)
GLM family binomial binomial binomial binomial poisson
Link logit cloglog.sub logit logit log
R formula I.sub~ C.sub I.sub~ C.sub Icheck.sub~ Ccheck.sub Itilde.sub~ Ctilde.sub, w=wtilde.sub N.sub~ C.sub
Offset term $\big{(}\log\frac{\pi_{1}}{\pi_{0}}$ $+\log\delta_{j}\big{)}_{j}$ $(\log\delta_{j})_{j}$ $\big{(}\log\frac{\pi_{1}}{\pi_{0}}$ $+\log\delta_{j}\big{)}_{j}$ $\big{(}\log\frac{\pi_{1}}{\pi_{0}}$ $+\log\delta_{j}\big{)}_{j}$ $\big{(}\log\big{(}\frac{\pi_{1}}{\pi_{0}}\big{)}$ $\mathbf{1}(j\in\mathcal{J}_{0})$ $+\log\left(\delta_{j}\right)\big{)}_{j}$
Size of the response vector #$\#\mathcal{J}_{0}+\#\mathcal{J}_{1}$$\#\mathcal{J}_{0}+\#\mathcal{J}_{1}$$\#\mathcal{N}+\#\mathcal{J}$$\#\mathcal{N}_{1}+\#\mathcal{J}$$\#\mathcal{J}_{0}+\#\mathcal{J}_{1}$
@@
Robustness $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
[…]
@@ 6.2 Results
Tables 3-4 summarize the results obtained. We report the values of prediction criteria PPV, AUC and WW detailed in Section 6.1, as well as the calculation time per bag. Implementation is realized within the R software (R Core Team, 2021) and in particular using the package glmnet (Friedman et al., 2010). Since the intensity function is estimated separately for each season and period of day, the computations are parallelized on a 20-core cluster. The overall computational time is essentially due to the estimation procedure while the cost of the prediction step mainly comes from the storage of predicted values for each season, period of day and pixel grid.
From Tables 3-4, a few comments can be formulated: first of all, whatever the criterion, it seems useless to choose more that one bag; second, note that AUC does not appear so informative, since all methods and subsampling rates achieve the same kind of performance. This is also the case for the PPV criterion in winter 00:00-06:00 (Table 4). Note also that PPV is quite small, in particular for winter (see Table 4). This confirms that the prediction of lightning storms is a very difficult task. However, this should be compared to the PPV achieved by the random classifier, which is the proportion of non-empty cells for these pairs of (season, moment) ($0.0154$ for (summer, 12:00-18:00) and $0.0004$ for (winter, 00:00-06:00)). Two methods ($\mathrm{PL}$ and $\mathrm{BRL}_{\text{\scriptsize\rm cloglog}}$) stand out for their stability concerning the subsampling rate, as demonstrated by the behaviors of PPV and WW in Table 3. Additionally, in the absence of subsampling ($\pi_{0}=1$), PPV lacks discrimination. However, in terms of WW, $\mathrm{PL}$ and $\mathrm{WCLRL}$ slightly outperform the other methods. Notably, subsampling significantly impacts computational time reduction (by at least a factor of 40) without compromising results.
@@ 7 Conclusion
We believe this work leads to several perspectives. It would be first interesting to study more deeply these procedures. For instance, can we still be ensured that the estimator derived by these procedures is consistent and asymptotically normal? Extending these procedures for non piecewise constant covariates. Investigating inhomogeneous subsampling, i.e. for instance letting $\pi_{0}$ and $\pi_{1}$ depend on the cell $\Delta_{j}$ are also interesting perspective. Finally, there is now a large literature on feature selection for point processes in particular using regularization techniques, e.g. Choiruddin et al. (2018). Understanding how these techniques can be combined with the zero-deflated subsampling procedure proposed in this paper is of particular relevance.
Regarding the dataset application, this paper has explored the spatio-temporal first-order trend of lightning strikes by assuming nothing on the second-order structure. This is standard when one anlyzes a point pattern, e.g. Waagepetersen and Guan (2009). Based on the present data analysis, modelling second-order characteristics of the process or even better modelling the distribution of $\mathbf{Y}$ for instance as a mixture of a temporal Hawkes process and a Neymann-Scott or log-Gaussian Cox process for the spatial part is a natural follow-up to the present work. Such a modelling could be in particular relevant if one intends to provide a short-term forecast of electrical activity.